### v1

@@ IV-C YCB Objects Perception
For this setting, we recorded at 60, 80, 100, 125, 150, 200, 250, and 300 cm from the cameras, where we used 10 different objects—see Fig. 12. The objects were selected to represent various shapes, sizes, and materials.
Fig. 12 includes front and side views of point clouds captured by the cameras at 60 cm and 300 cm. We can see that for the small distance, most of the objects are visually fine. Some flaws can be seen for can and windex bottle. The main difference between the cameras is for the pitcher, where the best results come from D435 and OAK-D Pro. D455 was unable to detect the hole in the handle, and ZED 2 missed half of the handle. The side views are worse again for the pitcher and then for the windex bottle. The bottle is transparent, which is an adversarial surface for RGB-D cameras. At 300 cm, the best front views are generally distorted and noisy and the best are probably for D455 and ZED 2. The more interesting is the side view, which reveals a higher standard deviation in the depth estimation from D435 and OAK-D Pro. Point clouds for D455 and ZED 2 have less scatter in depth estimation, but ZED 2 has noisy points and D455 creates surface coplanar with the table at the bottom parts of the objects.
@@
Chamfer Distance (CD) for the YCB objects can be seen in Fig. 13. We can see that the trend is the same as in the case in the plastic doll perception. Up to 150 cm all the cameras perform almost the same with CD around 1 cm. After that, the error for RealSense devices increases up to about 5.5 cm for D435. The best is ZED 2 with about 2 cm even at 300 cm. The difference from the plastic doll perception is that OAK-D Pro is now the second best camera in terms of mean CD. We assume that the reason is that some of the objects selected are flat, and thus OAK-D Pro performs well on these. But, as seen from a high variance and outlier points, it performs the worst on some objects.
@@
Angles between the closest normals for YCB objects are shown in Fig. 14. The results are similar to the angles for the plastic doll perception (see Fig. 9), with the difference for OAK-D Pro. In case of the plastic doll perception, there was a big difference between OAK-D Pro and the other cameras. Here, all cameras have similar results. It is again caused by the fact that some objects selected are flat, and therefore OAK-D Pro can estimate these well. However, the outliers (mainly at 60 cm) shows that it has problems with some objects.
@@
Jaccard Similarity (JS) for this scenario is shown in Fig. 15. We can see that at lower distances, the RealSense devices perform the best and ZED 2 getting better with increasing distance. At higher distances, the mean is similar for both ZED 2 and OAK-D Pro, but based on the variance and outliers, ZED 2 performs comparably on all objects. OAK-D Pro, on the other hand, has many outliers—objects for which the estimation is good, and thus the mean JS is higher. Analogous results are visible for the $F_{1}$ score in Fig. 16. At lower distances, all cameras achieve similar scores, but gradually ZED 2 is the best with the highest $F_{1}$ score and lower variance between objects.
@@
Fig. 17 provides per-object comparison of CD. We provide only CD, but all graphs can be found, together with the data used to create these, at https://osf.io/f2seb. Generally, we can say that the most problematic objects are a pitcher, a windex bottle, a football, and a drill. The windex bottle is transparent, which is a challenging surface for feature matching. As the bottle was on the table, the cameras could see the table through it, which probably aided perception. If there was nothing behind the bottle, the results would be even worse. The drill used has many small features on the surface that are hard to perceive, mainly at longer distances. In the case of football and pitcher, the main problem is probably the curvature—and, for the pitcher, the handle. We can see that the biggest error for these two have ZED 2 and OAK-D Pro. It supports the previous results and shows that these two cameras perform better with planar objects and suffer with more complex, curved objects.
YCB objects perception – summary. Overall, the results using YCB objects correspond to those using the plastic doll. All cameras are comparable up to 150 cm. After that, the best one is again ZED 2, now with OAK-D Pro being the second best. The reason why OAK-D Pro is not the worst as in the case of doll perception is that the dataset used contains some flat objects that are estimated correctly and decrease the overall error over all the objects.
@@ V Conclusion and Discussion
We compared four state-of-the-art stereoscopic depth cameras on more than 3,000 frames taken with each camera. The segmented data are available at https://osf.io/f2seb. Using six metrics, we compared the cameras on planar surface perception, plastic doll perception and on 10 household objects from the YCB dataset. The results show that even though the working principles are similar, individual cameras have different performance in different contexts and at different distances. In conclusion, on the basis of the results, we recommend the following. If your application requires an easy-to-setup and robust solution for lower distances (up to 90-100 cm) and is more object perception focused with unknown objects of arbitrary shapes, then the choice is RealSense D435. If you want to work at higher distances or also be able to detect planar surfaces properly (e.g., in mobile robots and for SLAM), select RealSense 455 or Stereolabs ZED 2. Here, the choice depends on whether you can afford to have a dedicated GPU for depth estimation (ZED 2) or not (D455). If you want to detect planar surfaces and/or objects that are less complex (e.g., boxes) the choice would be ZED 2 or Luxonis OAK-D Pro—again depending on whether you have a GPU (then choose ZED 2) or not (choose OAK-D Pro). In case you also require AI features, you can choose ZED 2 or OAK-D Pro. The choice here is based on whether you want to compute the AI on-board (OAK-D Pro) or have a dedicated computer GPU (ZED 2).
@@
The overall best camera is StereoLabs ZED 2. This camera provided the best results for planar surface estimation in terms of bias and standard deviation. Also, visually the point clouds created are the most flat. For plastic doll and YCB object perception, the camera also showed superior performance. For all metrics it achieved the best mean performance, while keeping the deviation between individual frames lower than the other cameras. It also provides easily accessible AI features like keypoint detection or face tracking. On the other hand, the camera requires CUDA-enabled GPU, which may be a drawback for some applications. ZED 2 also has a long baseline (the distance between two sensors in the physical camera) and works better from distances about 100 cm, which could be impractical for in some tasks like manipulation. On lower distances it tends to create holes in the object and miss some more complicated parts. From a user point of view, the API and associated software is not as intuitive as for Intel RealSense devices. At first, we had some problems setting the environment and installing the proper versions of the required libraries, which was never the case for RealSense cameras.
@@
The other cameras have a more specific use case. The Luxonis OAK-D Pro camera provides on-board AI features without the need for an external GPU. It works great for planar surface, where it provides performance comparable with ZED 2 camera, while computing everything on-board. However, it’s performance decreases a lot when perceiving more complex objects. The camera basically groups points in layers of constant depths. At lower distances, the layers are close to each other, so the overall perceived shape is correct. At higher distances, the layers are far from each other—see Fig. 5, Fig. 7, and Fig. 12. In addition, the camera seems to “load” when started, i.e., sometimes first frames after the start tend to be unfocused and more noisy. The API and software are the least intuitive of all the cameras.
@@
The RealSense D435 and D455 are the most typical RGB-D cameras currently used in robotics. They work on the same stereoscopic principle with IR projection computing everything on-board CPU without any additional AI features. The difference in these is the baseline that results in different performance. The D435 is smaller and therefore works at lower distances. The ideal range is only up to 3 m, so the performance decreases quickly with increasing distance, but up to 100 cm, the planar estimation is on par with other cameras (except for ZED 2, which is better) and for shape estimation it is overall the best in this range. The D455 seems to work better on the plane estimation task and works better for bigger distances. The RealSense API and all other software is the best among all the cameras. However, we also encountered some problems. Sometimes, the cameras would not start properly without removing and reinserting the USB cable. Also, mainly D455 seems to have a fixed order of turning on RGB and depth channels. If it is done in incorrect order, the depth and RGB images are not correctly matched. D455 also has a low RGB resolution (1280x800), which can be, mainly together with high FOV, a problem for RGB-based computer vision applications such as keypoint detection.
@@
Future work should also focus on comparing Time-of-Flight (ToF) sensors. Moreover, the cameras used are the state of the art in the time of publication, but the field is evolving rapidly, and it will be necessary to provide new studies when new sensors become available.

### v2

@@ IV-C YCB Objects Perception
For this setting, we recorded at 60, 80, 100, 125, 150, 200, 250, and 300 cm from the cameras, where we used 10 different objects—see Fig. 3 and Fig. 13. The objects were selected to represent various shapes, sizes, and materials.
Fig. 13 includes front and side views of point clouds captured by the cameras at 60 cm and 300 cm. We can see that for the small distance, most of the objects are visually fine. Some flaws can be seen for can and windex bottle. The main difference between the cameras is for the pitcher, where the best results come from the D435 and OAK-D Pro. D455 was unable to detect the hole in the handle, and ZED 2 missed half of the handle. The side views are worse again for the pitcher and then for the windex bottle. The bottle is transparent (see Fig. 3), which is an adversarial surface for RGB-D cameras. At 300 cm, the best front views are generally distorted and noisy and the best are probably for D455 and ZED 2. The more interesting is the side view, which reveals a higher standard deviation in the depth estimation from D435 and OAK-D Pro. Point clouds for D455 and ZED 2 have less scatter in depth estimation, but ZED 2 has noisy points and D455 creates surface coplanar with the table at the bottom parts of the objects.
@@
Chamfer Distance (CD) for the YCB objects can be seen in Fig. 14. We can see that the trend is the same as in the case in the plastic doll perception. Up to 150 cm all the cameras perform almost the same with CD around 1 cm. After that, the error for RealSense devices increases up to about 5.5 cm for D435. The best is ZED 2 with about 2 cm even at 300 cm. The difference from the plastic doll perception is that OAK-D Pro is now the second best camera in terms of mean CD. We assume that the reason is that some of the objects selected are flat, and thus OAK-D Pro performs well on these. But, as seen from high variance and outlier points, it performs the worst on some objects.
@@
Angles between the closest normals for YCB objects are shown in Fig. 15. The results are similar to the angles for the plastic doll perception (see Fig. 10), with the difference for OAK-D Pro. In case of plastic doll perception, there was a big difference between OAK-D Pro and the other cameras. Here, all cameras have similar results. This is again caused by the fact that some objects selected are flat, and therefore OAK-D Pro can estimate these well. However, the outliers (mainly at 60 cm) shows that it has problems with some objects.
@@
Jaccard Similarity (JS) for this scenario is shown in Fig. 16. We can see that at lower distances, the RealSense devices perform the best and ZED 2 is getting better with increasing distance. At higher distances, the mean is similar for both ZED 2 and OAK-D Pro, but based on the variance and outliers, ZED 2 performs comparably on all objects. OAK-D Pro, on the other hand, has many outliers—objects for which the estimation is good, and thus the mean JS is higher. Analogous results are visible for the $F_{1}$ score in Fig. 17. At lower distances, all cameras achieve similar scores, but gradually ZED 2 is the best with the highest $F_{1}$ score and lower variance between objects.
@@
Fig. 18 provides per-object comparison of CD. We provide only CD, but all graphs can be found, together with the data used to create them, at https://rustlluk.github.io/rgbd-comparison. Generally, we can say that the most problematic objects are a pitcher, a windex bottle, a football, and a drill—see Fig. 3 for RGB images of the objects. The windex bottle is transparent, which is a challenging surface for feature matching. As the bottle was on the table, the cameras could see the table through it, which probably aided perception. If there was nothing behind the bottle, the results would be even worse. The drill used has many small features on the surface that are difficult to perceive, mainly at longer distances. In the case of football and pitcher, the main problem is probably the curvature—and, for the pitcher, the handle. We can see that the biggest errors for these two have ZED 2 and OAK-D Pro. It supports the previous results and shows that these two cameras perform better with planar objects and suffer with more complex curved objects.
YCB objects perception – summary. Overall, the results using YCB objects correspond to those using the plastic doll. All cameras are comparable up to 150 cm. After that, the best one is again ZED 2, now with OAK-D Pro being the second best. The reason why OAK-D Pro is not the worst as in the case of doll perception is that the dataset used contains some flat objects that are estimated correctly and decrease the overall error over all the objects. This scenario is the most useful for table-top tasks, such as object perception or grasping. These tasks are usually performed up to a certain range from the robot, and therefore we would advise using D435 as it provides the best price/performance ratio for these tasks.
@@ V Conclusion and Discussion
We compared four state-of-the-art stereoscopic depth cameras on more than 3,000 frames taken with each camera. Segmented data are available at https://rustlluk.github.io/rgbd-comparison. Using six metrics, we compared the cameras on planar surface perception, plastic doll perception, and on 10 household objects from the YCB dataset. The results show that even though the working principles are similar, individual cameras have different performance in different contexts and at different distances. In conclusion, on the basis of the results, we recommend the following. If your application requires an easy-to-setup and robust solution for lower distances (up to 90-100 cm) and is more object perception focused with unknown objects of arbitrary shapes, then the choice is RealSense D435. If you want to work at higher distances or also be able to detect planar surfaces properly (e.g., in mobile robots and for SLAM), select RealSense D455 or Stereolabs ZED 2. Here, the choice depends on whether you can afford to have a dedicated GPU for depth estimation (ZED 2) or not (D455). If you want to detect planar surfaces and/or objects that are less complex (e.g. boxes), the choice would be ZED 2 or Luxonis OAK-D Pro—again depending on whether you have a GPU (then choose ZED 2) or not (choose OAK-D Pro). In case you also require AI features, you can choose ZED 2 or OAK-D Pro. The choice here is based on whether you want to compute the AI on-board (OAK-D Pro) or have a dedicated computer GPU (ZED 2).
It is important to look at all the metrics applicable in a given scenario. In case of planar surface perception, the metrics (bias and standard deviation) are the standard used in the literature [27, 33] and are, to the best of our knowledge, sufficient to properly evaluate the performance. However, this is not the case for the plastic doll and YCB objects perception, as none of the metrics is perfect and sufficient on its own. The Chamfer Distance (CD) is not robust to incorrect alignment of perceived and ground-truth point clouds, e.g., small rotation of one of the point clouds at the center of mass (mainly for bigger objects) results in high errors as closest points at the boundaries of the object are far away. Jaccard Similarity (JS) is robust to these as voxelization can mitigate the effect of small alignment errors. However, it is very sensitive to outliers, e.g., missing handle of a pitcher or hallucinated peaks. $F_{1}$ score is not informative in cases where there are no outliers. This can be seen in the case of the plastic doll, where the $F_{1}$ score (Fig. 12) is not changing until 175 cm, but for example the JS (Fig. 11) shows decreasing performance. The angle between normals allows to assess performance in the sense of shape similarity but does not capture the effect of outliers and missing points.
@@
The overall best camera is StereoLabs ZED 2. This is a similar finding to [33]. This camera provided the best results for planar surface estimation in terms of bias and standard deviation. Also, visually the point clouds created are the most flat. For plastic doll and YCB object perception, the camera also showed superior performance. For all metrics, it achieved the best mean performance, while keeping the deviation between individual frames lower than the other cameras. It also provides easily accessible AI features such as keypoint detection or face tracking. On the other hand, the camera requires CUDA-enabled GPU, which may be a drawback for some applications. ZED 2 also has a long baseline (the distance between two sensors in the physical camera) and works better from distances about 100 cm, which could be impractical for some tasks such as manipulation. On lower distances it tends to create holes in the object and miss some more complicated parts. From a user point of view, the API and associated software is not as intuitive as for Intel RealSense devices. At first, we had some problems setting the environment and installing the proper versions of the required libraries, which was never the case for RealSense cameras.
@@
The other cameras have a more specific use case. The Luxonis OAK-D Pro camera provides on-board AI features without the need for an external GPU. It works great for planar surface, where it provides performance comparable with ZED 2 camera, while computing everything on-board. However, it’s performance decreases a lot when perceiving more complex objects. The camera basically groups points into layers of constant depths. At lower distances, the layers are close to each other, so the overall perceived shape is correct. At higher distances, the layers are far from each other—see Fig. 6, Fig. 8, and Fig. 13. The same behavior was discovered in [33], where they discarded the camera from the comparison on a spherical object because the errors were too high. The effect is visible in the difference between JS and CD. Performance, for example in plastic doll perception with JS at higher distances, is comparable to the RealSense devices, but for CD the error is notably higher. We speculate that this behavior arises from optimizations in image resolution that allows all computations to be performed on board of the sensor. This makes the camera usable for more articulated objects only at lower distances. For simpler and more planar objects, it can be still used at higher distances and the onboard computations make it usable, for example, in indoor navigation and mapping. In addition, the camera seems to “load” when started, i.e., sometimes first frames after the start tend to be unfocused and more noisy. The API and software are the least intuitive of all the cameras—as also stated in [33].
@@
The RealSense D435 and D455 are the most typical RGB-D cameras currently used in robotics (this is supported by the fact that the camera are chosen in most of the state of the art in camera comparison, e.g. [33, 34, 27, 29]). They work on the same stereoscopic principle with IR projection computing everything using on-board CPU without any additional AI features. The difference in these is the baseline that results in different performance. The D435 has a smaller baseline and therefore works at lower distances. The ideal range is only up to 3 m, so the performance decreases quickly with increasing distance, but up to 100 cm, the planar estimation is on par with other cameras (except for ZED 2, which is better) and for shape estimation it is overall the best in this range. The D455 seems to work better on the plane estimation task and works better for larger distances—the same was discovered in [34, 33]. The stereo baseline is an important parameter, as with a fixed baseline the depth estimation error grows quadratically [41], which is nicely visible in Fig. 5. This is true for cameras with “classic” feature matching. We can see, again in Fig. 5, that it is not the case for ZED 2 which computes depth values using a neural network. The second parameter influenced by the baseline is the minimum distance in front of the camera—again, it is not true for ZED 2. The bigger the baseline, the larger the necessary distance in front of the camera. This property is a problem mainly for close proximity sensing—see [13].
The RealSense API and all other software is the best among all the cameras. However, we also encountered some problems. Sometimes, the cameras would not start properly without removing and reinserting the USB cable. Also, mainly D455 seems to have a fixed order of turning on RGB and depth channels. If it is done in incorrect order, the depth and RGB images are not correctly matched. D455 also has a low RGB resolution (1280x800), which can be, mainly together with high FOV, a problem for RGB-based computer vision applications such as keypoint detection.
Light conditions during sensing are important, which is also tied to the fact whether the cameras use active (IR projection) or passive technology. et al.Halmetschlager-Funek [27] showed that light intensity influences the results and IR projection can help in darker environments and also with reflective surfaces in brighter environments. Furthermore, in [33] the authors showed that ZED 2, as a passive camera without IR projection, has shown lower performance under low lighting. Also, when there is direct sunglight, performance is very limited [30]. We tried to keep the light conditions as consistent as possible. However, we were unable to change them to test different ones. Such experiments could be added in the future.
@@
Future work should also focus on comparing Time-of-Flight (ToF) sensors. Moreover, the cameras used are the state of the art at the time of publication, but the field is evolving rapidly, and new studies will be necessary when new sensors become available.